OpenStudy (anonymous):

How much heat is required to raise the temperature of 25 grams of water from 25.0 degrees Celsius to 30.0 degrees Celsius? The specific heat of water is 4.18 J/ g multiplied by degrees Celsius.

OpenStudy (anonymous):

\[\rm Heat = mass \times specific ~ heat \times change ~ i n ~ temperature\]\[Q = m \times c \times \Delta T\]

OpenStudy (anonymous):

thanks so much...it really helps

OpenStudy (anonymous):

would 522.5 be the answer?

OpenStudy (anonymous):

Yep!